OpenStudy (anonymous):

A 3,000-N force gives an object an acceleration of 15 m/s². The mass of the object is____.

OpenStudy (anonymous):

f = ma; a = m/f; m =f/a where f = force m = mass in kg a = acceleration m = 3000N/15 m/s² = 200 kg

OpenStudy (anonymous):

***a = f/m

OpenStudy (anonymous):

thanks Tutor_guy you're best